# Gravitational-wave signatures of non-violent non-locality

### v2

@@ Abstract
We begin with the effective-one-body (EOB) framework [37, 38, 39] which approximates a two-body relativistic problem as a one-body problem in a deformed Schwarzschild spacetime, the effective spacetime. Originally, this was developed for nonspinning quasicircular [37, 38, 39, 40, 41, 42] but subsequently was generalized to aligned spin [43, 44, 45, 46], precessing spin [47, 48, 49, 50, 51, 52], calibrated to NR [53, 54, 55], tides [56, 57, 58, 59], and eccentricity [60, 61] where the newest model is SEOBNRv5PHM [62, 63, 64, 65]. We modify a simple nonspinning EOBNRv2 spacetime [40] and add stochastic metric perturbations characterized by a Gaussian spatial profile and a frequency spectrum related to the BH’s temperature, as proposed by Ref. [21]. The resulting trajectories lead to GW that deviate from GR stochastically, primarily in the late-inspiral and plunge phases. Using a principal component analysis (PCA) [66, 67, 68], the frequency-domain phase deviations of NVNL waveforms are well approximated by a single dominant eigenmode multiplied by a normally distributed random amplitude. Finally, we estimate how well NVNL can be constrained by stacking together gravitational-wave events and computing the Savage-Dickey ratio [69]. These calculations give theoretical support to search for random phase deviations as proposed in the hierarchical tests of GR [35, 29, 70, 71, 72, 73].
@@
[…]
For a test particle falling into a NVNL BH, using $\epsilon$ and $\eta$ as expansion parameters in particle mass and size of NVNL, we can perform a two-parameter expansion for the metric as $g=g^{\mathrm{S}}+\eta n+\epsilon h^{(0,1)}+\eta\epsilon h^{(1,1)}$, and the trajectory of the as $x^{(0,0)}+\eta x^{(1,0)}$. The modified Einstein equation, up to $O(1,1)$ can be written as $G[g+\eta n]=8\pi\eta T_{\rm NVNL}$ and
@@
where $G^{(1)}_{g+\eta n}$ denotes linearized Einstein tensor with background $g+\eta n$ and $T_{g+\eta n}[x]$ denotes stress-energy tensor of particle in this background with world line $x$. From Eq. (4), we recover $G^{(1)}_{g}[h^{(0,1)}]=8\pi T_{g}[x^{(0,0)}]$, obtain that $x^{(0,0)}+\eta x^{(1,0)}$ is a geodesic in $g+\eta n$ (linear Bianchi identity), and that
@@
Equation (4) indicates that GWs up to $O(1,1)$ is generated by a modified trajectory and propagate on the modified spacetime $g+\eta n$; (5) shows further that the propagation effect in $h^{(1,1)}$ can be computed as being sourced by a beat between $h^{(0,1)}$ and the metric perturbation $n$. Within GR, radiation reaction effect shows up in $x^{(0,1)}$ and $h^{(0,2)}$, while NVNL corrects in $x^{(1,1)}$ and $h^{(1,2)}$. In this paper, we shall restrict ourselves to the test-particle limit since, as we see in the Supplemental Materials, the NVNL is confined near the end of the inspiral and the finite-mass effects will not cumulate secularly. We will further neglect the wave propagation effect, since that involves the scattering of the merger wave and causes a modified ringdown wave — while we will be dealing with inspiral tests for NVNL.
@@
NVNL Waveforms. In Fig. 1, we plot several realizations of an NVNL waveform $\mathrm{Re}[h_{22}]$ (normalized by $M/D$ with $D$ the source distance) in the time domain, as functions of $t/M$, for a binary with $q=1$, and compare this to a GR waveform. The random deviations are smooth in time due to the cutoff $f_{Q}$ in Eq. (3) as well as the filtering effect due to the inertia of the binary.
@@
where we are using the notation $p(x|d,M)$ to represent the posterior probability density $x$ given data $d$ under modeling assumptions $M$. One can see that the Savage-Dickey ratio compares how much the posterior has changed to the prior at the location of the GR limit. We use priors which are uniform in the range of $-1\leq\mu\leq 1$ and $0\leq\sigma\leq 1$.
Next, we investigate the detectability of NVNL via signal injection. In Fig. 4 we show a contour plot of the Savage-Dickey ratio for a three detector network Livingston-Hanford-Virgo at O3 Livingston sensitivity. In particular, we calculate the Fisher information matrix for five years of events where we draw from the astrophysical rates. The merger rate density scales with the star formation rate and the masses are drawn from the Power Law+Peak (PP) model [83, 84] that is the best fit point from GWTC3 data [85]. We show the estimated measurement precision in the astrophysical parameters, marginalizing over the event-level parameters. We perform this analysis for the optimal PCA dephasing term. Additionally, we give the results for the traditional PN deformation coefficients based upon the biased framework [80]. We can see that NVNL can be constrained such that $A\lesssim 1.6\times 10^{-2}$ after five years of observation at O3 sensitivities when using the PCA method $\zeta$. The 3.5PN dephasing term, $\delta\varphi_{7}$, performed the best out of the parameterized tests and found the constraint $A\lesssim 2.1\times 10^{-2}$. We note that for a single BBH event $A\lesssim 0.1$ which is comparable what [86] found using a different scheme (c.f. Table 1 [86]). The minimum detectable value depend on weakly on PN order, but the PCA method was optimal. This is consistent with previous work showing that deviations are detectible with most PN tests [87, 88]. For third-generation detectors, we consider three Cosmic Explorer detectors located at the locations of the current LIGO/Virgo network. The five-year constraint on $A$ is now $A\lesssim 1.2\times 10^{-3}$ for the PCA method and $A\lesssim 1.7\times 10^{-3}$ for the 3.5PN parameterized test, as shown in Fig. 5. For Figs. 4 and 4, we resampled many realizations of events and their associated noise realization so that very loud events do not show drastic shifts in the constraint.
@@
$\frac{\partial p_{\phi}}{\partial t}$ =^$=\hat{\mathcal{F}}_{\phi}-\frac{\partial\left(h_{vv}^{\ell m}\right)}{\partial\phi}\Delta\hat{H}_{\ell m}^{\mathrm{real}}\,,$ (26)
@@
^$\hat{\mathcal{F}}_{\phi}=-\frac{1}{\eta v_{\omega}^{3}}\frac{dE}{dt}\,,$ (27)
@@
Figure 7: The log Bayes factor for the Hanford-Livingston-Virgo network at O3 Livingston sensitivity (positive favors GR). One can see that the red curve favors NVNL at large events, while blue favors GR. The whiskers correspond to the upper and lower bounds are ±$\pm 1\sigma$ percentile values due to randomness associated with the order of events. If one sees a loud clear event early, then it is easier to favor/disfavor GR.
Variance of Bayes Factor. As we noted in the discussion of Fig. 4 and Fig. 5, the Bayes factor measurement depends on randomness about the event order. In Fig. 7, we plot how the (log) Bayes Factor ratio scales for multiple injection sizes of $A$. In the center, we show the median Bayes factor for each injection size while the upper and lower bounds are ±$\pm 1\sigma$ percentile values for the Bayes factor after this many observations. This mostly occurs because the loudest events are the most informative, so the order of events can affect the rolling constraint. Furthermore, the statistical realization of the detector noise and hierarchical model draw add subdominant variations to the Bayes factor.

### v3

@@ Abstract
We begin with the effective-one-body (EOB) framework [37, 38, 39] which approximates a two-body relativistic problem as a one-body problem in a deformed Schwarzschild spacetime, the effective spacetime. Originally, this was developed for nonspinning quasicircular [37, 38, 39, 40, 41, 42] but subsequently was generalized to aligned spin [43, 44, 45, 46], precessing spin [47, 48, 49, 50, 51, 52], calibrated to NR [53, 54, 55], tides [56, 57, 58, 59], and eccentricity [60, 61] where the newest model is SEOBNRv5PHM [62, 63, 64, 65]. We modify a simple nonspinning EOBNRv2 spacetime [40] and add stochastic metric perturbations characterized by a Gaussian spatial profile and a frequency spectrum related to the BH’s temperature, as proposed by Ref. [21]. The resulting trajectories lead to GWs that deviate from GR stochastically, primarily in the late-inspiral and plunge phases. Using a principal component analysis (PCA) [66, 67, 68], the frequency-domain phase deviations of NVNL waveforms are well approximated by a single dominant eigenmode multiplied by a normally distributed random amplitude. Finally, we estimate how well NVNL can be constrained by stacking together gravitational-wave events and computing the Savage-Dickey ratio [69]. These calculations give theoretical support to search for random phase deviations as proposed in the hierarchical tests of GR [35, 29, 70, 71, 72, 73].
@@
where $f_{Q}=1/8\pi M$ is the quantum frequency scale. We have normalized the power spectrum so that $n(t)$ has a variance equal to unity $\langle n^{2}(t)\rangle\equiv\int dfS_{n}(f)=1$. This spectrum is motivated by the Boltzmann distribution for a BH of temperature $T_{\mathrm{BH}}=1/8\pi M$. Note that the standard deviation of $f_{\ell m}$ at the horizon is $\sim A_{\ell m}$, while the coherence time is around $\tau\sim 4M$.
For a test particle falling into a NVNL BH, using $\epsilon$ and $\eta$ as expansion parameters in particle mass and size of NVNL, we can perform a two-parameter expansion for the metric as $g=g^{\mathrm{S}}+\eta n+\epsilon h^{(0,1)}+\eta\epsilon h^{(1,1)}$, and the trajectory of the particle as $x^{(0,0)}+\eta x^{(1,0)}$. The modified Einstein equation, up to $O(1,1)$ can be written as $G[g+\eta n]=8\pi\eta T_{\rm NVNL}$ and
@@
where $G^{(1)}_{g+\eta n}$ denotes the linearized Einstein tensor with background $g+\eta n$ and $T_{g+\eta n}[x]$ denotes stress-energy tensor of a particle in this background with world line $x$. From Eq. (4), we recover $G^{(1)}_{g}[h^{(0,1)}]=8\pi T_{g}[x^{(0,0)}]$, obtain that $x^{(0,0)}+\eta x^{(1,0)}$ is a geodesic in $g+\eta n$ (linear Bianchi identity), and that
@@
Equation (4) indicates that GWs up to $O(1,1)$ are generated by a modified trajectory and propagate on the modified spacetime $g+\eta n$; (5) shows further that the propagation effect in $h^{(1,1)}$ can be computed as being sourced by a beat between $h^{(0,1)}$ and the metric perturbation $n$. Within GR, radiation reaction effects show up in $x^{(0,1)}$ and $h^{(0,2)}$, while NVNL corrects in $x^{(1,1)}$ and $h^{(1,2)}$. In this paper, we shall restrict ourselves to the test-particle limit since, as we see in the Supplemental Materials, the NVNL is confined near the end of the inspiral and the finite-mass effects will not cumulate secularly. We will further neglect the wave propagation effect, since that involves the scattering of the merger wave and causes a modified ringdown wave — while we will be dealing with inspiral tests for NVNL.
@@
NVNL Waveforms. In Fig. 1, we plot several realizations of an NVNL waveform $\mathrm{Re}[h_{22}]$ (normalized by $M/D$ with $D$ the source distance) in the time domain, as functions of $t/M$, for a binary with $q=1$, and compare this to a GR waveform. The random deviations are smooth in time due to the cutoff $f_{Q}$ in Eq. (3) as well as the filtering effect due to the inertia of the binary which suppresses high frequencies.
@@
where we are using the notation $p(x|d,M)$ to represent the posterior probability density $x$ given data $d$ under modeling assumptions $M$. One can see that the Savage-Dickey ratio compares how much the posterior has changed to the prior at the location of the GR limit. We use priors which are uniform in the range of $1\leq\mu\leq 1$ and $0\leq\sigma\leq 1$.
Next, we investigate the detectability of NVNL via signal injection. In Fig. 4 we show a contour plot of the Savage-Dickey ratio for a three detector network Livingston-Hanford-Virgo at O3 Livingston sensitivity. In particular, we calculate the Fisher information matrix for five years of events where we draw from the astrophysical rates. The merger rate density scales with the star formation rate and the masses are drawn from the Power Law+Peak (PP) model [83, 84] that is the best fit point from GWTC3 data [85]. We show the estimated measurement precision in the astrophysical parameters, marginalizing over the event-level parameters. We perform this analysis for the optimal PCA dephasing term. Additionally, we give the results for the traditional PN deformation coefficients based upon the biased framework [80]. We can see that NVNL can be constrained such that $A\lesssim 1.6\times 10^{-2}$ after five years of observation at O3 sensitivities when using the PCA method $\zeta$. The 3.5PN dephasing term, $\delta\varphi_{7}$, performed the best out of the parameterized tests and found the constraint $A\lesssim 2.1\times 10^{-2}$. We note that for a single BBH event $A\lesssim 0.1$ which is comparable what [86] found using a different scheme (c.f. Table 1 [86]). The minimum detectable value depends on weakly on PN order, but the PCA method was optimal. This is consistent with previous work showing that deviations are detectible with most PN tests [87, 88]. For third-generation detectors, we consider three Cosmic Explorer detectors located at the locations of the current LIGO/Virgo network. The five-year constraint on $A$ is now $A\lesssim 1.2\times 10^{-3}$ for the PCA method and $A\lesssim 1.7\times 10^{-3}$ for the 3.5PN parameterized test, as shown in Fig. 5. For Figs. 4 and 4, we resampled many realizations of events and their associated noise realization so that very loud events do not show drastic shifts in the constraint.
@@
$\frac{\partial p_{\phi}}{\partial t}$ $=\hat{\mathcal{F}}_{\phi}-\frac{\partial\left(h_{vv}^{\ell m}\right)}{\partial\phi}\Delta\hat{H}_{\ell m}^{\mathrm{real}}\,$, (26)
@@
$\hat{\mathcal{F}}_{\phi}=-\frac{1}{\eta v_{\omega}^{3}}\frac{dE}{dt}\,$, (27)
@@
Figure 7: The log Bayes factor for the Hanford-Livingston-Virgo network at O3 Livingston sensitivity (positive favors GR). One can see that the red curve favors NVNL at large events, while blue favors GR. The whiskers correspond to the upper and lower bounds are $\pm 1\sigma$ percentile values due to randomness associated with the order of events. If one sees a loud clear event early, then it is easier to favor/disfavor GR.
Variance of Bayes Factor. As we noted in the discussion of Fig. 4 and Fig. 5, the Bayes factor measurement depends on randomness about the event order. In Fig. 7, we plot how the (log) Bayes Factor ratio scales for multiple injection sizes of $A$. In the center, we show the median Bayes factor for each injection size while the upper and lower bounds are $\pm 1\sigma$ percentile values for the Bayes factor after this many observations. This mostly occurs because the loudest events are the most informative, so the order of events can affect the rolling constraint. Furthermore, the statistical realization of the detector noise and hierarchical model draw add subdominant variations to the Bayes factor.